OpenStudy (anonymous):

Anyone taking Connections Academy Geometry Semester 1??? I need help!

OpenStudy (anonymous):

What's your question

OpenStudy (anonymous):

Help with what?

OpenStudy (anonymous):

@MariaSmith10 I won't give answers, but I can definitely help you through the problems.

OpenStudy (anonymous):

@ me when you have a question @MariaSmith10

OpenStudy (anonymous):

Is the line through the points R(-1, 3) and S(2, -7) parallel to the graph of the line given by the equation, 10x + 3y = 6?

OpenStudy (anonymous):

Change 10x+3y = 6 to slope intercept form y = mx+b 3y = 6 - 10x y = 2 - 10/3x The slope is -10/3 so the line we're going to look at needs the same slope

OpenStudy (anonymous):

slope = m = y2-y1 / x2-x1 -7-3 /2 + 1 = -10/3 Same slope as before to the lines are parallel

OpenStudy (anonymous):

a. No, the slopes have opposite signs. b. Yes, the slopes have the same sign and value. c. Yes, the lines both decrease to the right. d. No, the lines have unequal slopes.

OpenStudy (anonymous):

I would say B...both slopes have the same sign and value...if you were to graph, they would also decrease from left to right. Can you have more than one answer

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I'd say B is the best answer

OpenStudy (anonymous):

ok thanks.